OpenStudy (anonymous):

Integral of e^(x^(1/3))? I'd imagine some combination of integration by parts and u-substitution are required, but it's not apparent at the moment. An explanation would be ten times more valuable than just an answer.

zepdrix (zepdrix):

Hmm let's try a u-sub and see if it will work...

OpenStudy (anonymous):

with what, though? The obvious thing to try would be u=cube root of x, but we don't have a du in there.

zepdrix (zepdrix):

\[\Large\bf\sf \int e^{x^{1/3}}\;dx\] \[\Large\bf\sf u=x^{1/3}\]\[\Large\bf\sf du=\frac{1}{3}x^{-2/3}\;dx\qquad\to\qquad 3x^{2/3}du=dx\]Which we can write as,\[\Large\bf\sf 3(x^{1/3})^2du=dx\]\[\Large\bf\sf 3u^2\;du =dx\]

zepdrix (zepdrix):

We have to uhhh, "make" a du. It's a little tricky.

OpenStudy (anonymous):

OK! That makes sense!

zepdrix (zepdrix):

Yah I'll bet parts works nicely from there :)

OpenStudy (anonymous):

so now we have \[3u^2e^udu\]

OpenStudy (anonymous):

which works pretty well with integration by parts, I think.

OpenStudy (anonymous):

\[\int\limits3u^2e^udu = 3u^2e^u-6ue^u+6e^u\]

OpenStudy (anonymous):

Thanks! I'll remember to try that trick next time I'm stuck - is there a formal name for it?

zepdrix (zepdrix):

Hmm I'm not sure! It's just a neat little tricky where you sometimes have to use your `u` within your `du`.